Direct Materials Purchases Budget Romanos Frozen Pizza Inc. has determined from its production budget the following estimateDough Tomato Cheese Total Units required for production: 12 pizza 16 pizza Desired inventory, September 30 Total units avai
Answer #1
Romano's Frozen Pizza Inc.
Direct Materials Purchases Budget
For the Month Ending September 30
Dough Tomato Cheese Total
Units required for production:
12" pizza 14490 9660 12880
16" pizza 37200 24800 32240
Desired inventory, September 30 550 230 350
Total units available 52240 34690 45470
Estimated inventory, September 1 -520 -240 -320
Total units to be purchased 51720 34450 45150
Unit price 1.30 2.40 3.40
Total direct materials to be purchased 67236 82680 153510 303426
Workings:
Dough Tomato Cheese
Units required for production:
12" pizza =16100*0.9 =16100*0.6 =16100*0.8
16" pizza =24800*1.5 =24800*1 =24800*1.3